# Sample Complexity of Linear Quadratic Regulator Without Initial Stability

### v2

@@ 3. The Outer Loop (Receding-Horizon Policy Gradient)
It has been demonstrated that the solution to the RDE (20) converges monotonically to the stabilizing solution of the ARE (4) exponentially [7]. As a result, $\{K^{*}_{t}\}_{t\in\{N-1,\dotsc,1,0\}}$ in (21) also converges monotonically to $K^{*}$ as $N$ increases. In particular, we recall the following result from [20, Theorem 1].
@@
The proof of Theorem 3.1 is provided in Appendix A for completeness (and to account for some minor change in notation). We also note that this theorem relies on a minor inherent assumption that $Q_{N}$ satisfies $Q_{N}\succeq P^{*}$. A full discussion of this assumption is provided in Remark A.1 in Appendix A.
With this result in place, we provide our proposed algorithm (see Algorithm 1). Note that in this section, we focus on the outer loop of Algorithm 1, analyzing the requirements it imposes on the convergence of the policy gradient method employed in the inner loop. The details of the policy gradient method will be discussed in the next section.
@@
0:  Horizon $N$, max iterations $\{T_{h}\}$, stepsizes $\{\alpha_{h,t}\}$, variance parameter $\sigma^{2}$
1:  for $h=N-1$ to 00 do
2:     Initialize $K_{h,0}$ arbitrarily (e.g., the convergent policy from the prev. iter. $K_{h+1,T_{h+1}}$ or 00).
@@
4:        Sample $x_{h}\sim\mathcal{D}$ and $\eta_{h,t}\sim\mathcal{N}(0,I_{m})$ and simulate a trajectory with $u_{h,t}=-K_{h,t}x_{h}+\sigma\eta_{h,t}$.
5:        Compute $Q_{h}(x_{h},u_{h,t})$ for said trajectory.
@@
^$\widehat{\nabla}\widetilde{J}_{h,t}(K_{h,t})=-\frac{1}{\sigma}Q_{h}(x_{h},u_{h,t})\eta_{h}x_{h}^{\top}.$
7:        Update $K_{h,t+1}=K_{h,t}-\alpha_{h,t}\cdot\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$.
@@
$a:=\frac{\sigma_{\min}(Q)}{2}$
$\varphi:=\max_{t\in\{0,1,\dotsc,N-1\}}\|A-BK^{*}_{t}\|$
$P_{\max}:=\max_{t\in\{0,1,\dotsc,N-1\}}\{P^{*}_{t}\}$
$C_{1}:=\frac{\varphi\|B\|}{\lambda_{\min}(R)}$
$C_{2}:=2\varphi\|A\|\left(1+\frac{\|P_{\max}+aI\|\|B\|^{2}}{\lambda_{\min}(R)}\right)$
$C_{3}:=2\|R+B^{\top}(P_{\max}+aI)B\|.$
@@
(29) $\varsigma_{h,\varepsilon}:=\begin{cases}\min\left\{\sqrt{\frac{a}{C_{3}N}},\sqrt{\frac{a^{2}}{2eC_{3}N\|P_{\max}\|}},\sqrt{\frac{a\varepsilon}{8eC_{3}NC_{1}\|P_{\max}\|}},\sqrt{\frac{\varepsilon}{4C_{1}C_{3}}}\right\},&h\geq 1,\\ \frac{\epsilon}{4},&h=0.\end{cases}$
We now present a key result, Theorem 3.2, on the accumulation of errors that constitutes an improvement over [20, Theorem 2] (corrected version of which is stated as Theorem 3.3 below); as the proof of Theorem 3.2 demonstrates, this improvement relies on a fundamentally different analysis.
@@
(30) $N=\frac{1}{2}\cdot\frac{\log\left(\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\epsilon\cdot\lambda_{\min}(R)}\right)}{\log\left(\frac{1}{\|A_{K}^{*}\|_{*}}\right)}+1,$
where $Q_{N}\succeq P^{*}$, and suppose that Assumption 2.1 holds. Now assume that, for some $\epsilon>0$, there exists a sequence of policies $\{\widetilde{K}_{h}\}_{h\in\{N-1,\dotsc,0\}}$ such that for all $h\in\{N-1,\dotsc,0\}$,
@@
where $\widetilde{K}^{*}_{h}$ is the optimal policy for the Linear Quadratic Regulator (LQR) problem from step $h$ to $N$, incorporating errors from previous iterations of Algorithm 1. Then, the proposed algorithm outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$. Furthermore, if $\varepsilon$ is sufficiently small such that
$\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\|B\|},$
then $\widetilde{K}_{0}$ is stabilizing.
The proof of Theorem 3.2 is presented in Appendix B. A key component of our analysis is the contraction of the Riemannian distance on the Riccati operator, as established in Lemma 2.1. This allows us to demonstrate that the accumulated error remains linear in $N$, in contrast to the exponential growth in [20, Theorem 2].
Given this discrepancy, we revisit [20, Theorem 2] and present a revised version which accounts for some necessary, and non-trivial, modifications to make the statement accurate. For the latter reason, and the fact that this result does not rely on Assumption 2.1, we provide a complete proof in Appendix C.
@@
(31) $N=\frac{1}{2}\cdot\frac{\log\left(\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\epsilon\cdot\lambda_{\min}(R)}\right)}{\log\left(\frac{1}{\|A_{K}^{*}\|_{*}}\right)}+1,$
where $Q_{N}\succeq P^{*}$. Now assume that, for some $\epsilon>0$, there exists a sequence of policies $\{\widetilde{K}_{h}\}_{h\in\{N-1,\dotsc,0\}}$ such that
(32) $\|\widetilde{K}_{h}-\widetilde{K}^{*}_{h}\|\leq\begin{cases}\min\left\{\sqrt{\frac{a}{C_{3}}},\sqrt{\frac{a}{C_{2}^{h-2}C_{3}}},\frac{1}{2}\sqrt{\frac{\epsilon}{C_{1}C_{2}^{h-2}C_{3}}}\right\},&h\geq 2,\\[6.0pt] \min\left\{\sqrt{\frac{a}{C_{3}}},\frac{1}{2}\sqrt{\frac{\epsilon}{C_{1}C_{3}}}\right\},&h=1,\\[6.0pt] \frac{\epsilon}{4},&h=0.\end{cases}$
where $\widetilde{K}^{*}_{h}$ is the optimal policy for the Linear Quadratic Regulator (LQR) problem from step $h$ to $N$, incorporating errors from previous iterations of Algorithm 1. Then, the RHPG algorithm outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$. Furthermore, if $\varepsilon$ is sufficiently small such that
$\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\|B\|},$
then $\widetilde{K}_{0}$ is stabilizing.
@@ 4. The Inner Loop and Policy Gradient
We seek a way to estimate the gradient of this function with respect to $K_{h}$. To remedy, we propose:
(33) ^$\widehat{\nabla}\widetilde{J}_{h}(K):=-\frac{1}{\sigma^{2}}Q_{h}(x_{h},u_{h})(u_{h}+Kx_{h})x_{h}^{\top},$
where $x_{h}$ is sampled from $\mathcal{D}$, and then $u_{h}$ is chosen randomly from the Gaussian distribution $\mathcal{N}(-Kx_{h},\sigma^{2}I_{m})$ for some $\sigma>0$. Moreover, we rewrite $u_{h}\sim\mathcal{N}(-Kx_{h},\sigma^{2}I_{m})$ as
(34) $u_{h}=-Kx_{h}+\sigma\eta_{h},$
where $\eta_{h}\sim\mathcal{N}(0,I_{m})$. Substituting (34) in (33) yields
(35) ^$\widehat{\nabla}\widetilde{J}_{h}(K)=-\frac{1}{\sigma}Q_{h}(x_{h},-Kx_{h}+\sigma\eta_{h})\eta_{h}x_{h}^{\top}.$
@@
Suppose $x_{h}$ is sampled from $\mathcal{D}$ and $u_{h}$ chosen from $\mathcal{N}(-Kx_{h},\sigma^{2}I_{m})$ as before. Then for any given choice of $K$, we have that
@@
${\mathbb{E}}_{x_{h}}[\widehat{\nabla}\widetilde{J}_{h}(K)]$ $={\mathbb{E}}_{x_{h}}\left[{\mathbb{E}}_{\eta_{h}}\left[\widehat{\nabla}\widetilde{J}_{h}(K)\big{|}x_{h}\right]\right]$
=$\overset{\mathrm{(i)}}{=}{\mathbb{E}}_{x_{h}}\left[-\frac{1}{\sigma^{2}}{\mathbb{E}}_{\eta_{h}}\left[Q(x_{h},-Kx_{h}+\sigma\eta_{h})(\sigma\eta_{h})\big{|}x_{h}\right]x_{h}^{\top}\right]$
(37) =$\overset{\mathrm{(ii)}}{=}{\mathbb{E}}_{x_{h}}\left[{\mathbb{E}}_{\eta_{h}}\left[-\nabla_{u}Q^{K}(x_{h},u)\bigg{|}_{u=-Kx_{h}+\sigma\eta_{h}}\big{|}x_{h}\right]x_{h}^{\top}\right],$
where (i) follows from $x_{h}^{\top}$ being determined when given $x_{h}$, and (ii) from Stein’s lemma [15]. Using (15), we compute
$\nabla_{u}Q_{h}(x_{h},u)$ $=\nabla_{u}\Bigg{(}x_{h}^{\top}Qx_{h}+u^{\top}Ru+(Ax_{h}+Bu)^{\top}\widetilde{P}_{h+1}(Ax_{h}+Bu)\Bigg{)}$
$=2Ru+2B^{\top}\widetilde{P}_{h+1}Bu+2B^{\top}\widetilde{P}_{h+1}Ax_{h},$
which evaluated at $u=-Kx_{h}+\sigma\eta_{h}$ yields
$\nabla_{u}Q_{h}(x_{h},u)\bigg{|}_{u=-Kx_{h}+\sigma\eta_{h}}=2\left((R+B^{\top}\widetilde{P}_{h+1}B)(-Kx_{h}+\sigma\eta_{h})+B^{\top}\widetilde{P}_{h+1}Ax_{h}\right).$
@@
${\mathbb{E}}[\widehat{\nabla}\widetilde{J}_{h}(K)]$ $={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[2\left((R+B^{\top}\widetilde{P}_{h+1}B)K-B^{\top}\widetilde{P}_{h+1}A\right)x_{h}x_{h}^{\top}\right]$
$=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K-B^{\top}\widetilde{P}_{h+1}A\right){\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[x_{h}x_{h}^{\top}\right]$
@@
=$\overset{\mathrm{(i)}}{=}\nabla\widetilde{J}_{h}(K),$
@@
We now provide some bounds in the following lemma.
@@
Suppose $\zeta\in(0,\frac{1}{e}]$, and
$\|\widetilde{P}_{h+1}-P^{*}_{h+1}\|\leq a.$
Then for any $K\in{\mathcal{G}}_{h}$, we have that
(39) ^$\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}\leq\xi_{h,3}\left(\log\frac{1}{\zeta}\right)^{3/2}$
with probability at least $1-\zeta$, where $\xi_{h,1},\xi_{h,2},\xi_{h,3}\in{\mathbb{R}}$ are given by
(40) $\xi_{h,1}$ $:=\Big{(}\|Q\|+2\|R\|\widetilde{C}_{h}^{2}+2(\|P_{\max}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h})\Big{)}C_{m}^{3/2},$
(41) $\xi_{h,2}$ $:=2\left(\|R\|+2(\|P_{\max}\|+a)\|B\|^{2}\right)C_{m}^{1/2},$
(42) $\xi_{h,3}$ $:=\frac{1}{\sigma}\left(\xi_{h,1}5^{1/2}m^{1/2}\right)+\sigma\left(\xi_{h,2}5^{3/2}m^{3/2}\right).$
Moreover,
(43) ${\mathbb{E}}\left[\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}^{2}\right]\leq\xi_{h,4},$
where
(44) $\xi_{h,4}$ $:=\frac{1}{\sigma^{2}}\xi_{h,1}^{2}m+2\xi_{h,1}\xi_{h,2}m(m+2)+\sigma^{2}\xi_{h,2}^{2}m(m+2)(m+4).$
@@
Using the Formulation of ^$\widehat{\nabla}\widetilde{J}_{h}(K)$ derived in (35), we have
(45) ^$\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}$ $=\|\frac{1}{\sigma}Q_{h}(x_{h},-Kx_{h}+\sigma\eta_{h})\eta_{h}x_{h}^{\top}\|_{F}$
(46) $\leq\frac{1}{\sigma}Q_{h}(x_{h},-Kx_{h}+\sigma\eta_{h})\|\eta_{h}\|\|x_{h}\|.$
@@
(48) $\widetilde{J}_{h}(K)$ $\geq{\mathbb{E}}_{x_{h}\sim{\mathcal{D}}}\left[x_{h}^{\top}(Q+K^{\top}RK)x_{h}\right]$
(49) $={\mathbb{E}}_{x_{h}\sim{\mathcal{D}}}\left[\operatorname{tr}\left((Q+K^{\top}RK)x_{h}x_{h}^{\top}\right)\right]$

### v3

@@ 3. The Outer Loop (Receding-Horizon Policy Gradient)
It has been demonstrated that the solution to the RDE (20) converges monotonically to the stabilizing solution of the ARE (4) exponentially [9]. As a result, $\{K^{*}_{t}\}_{t\in\{N-1,\dotsc,1,0\}}$ in (21) also converges monotonically to $K^{*}$ as $N$ increases. In particular, we recall the following result from [22, Theorem 1].
@@
The proof of Theorem 3.1 is provided in Appendix A for completeness (and to account for some minor change in notation). We also note that this theorem relies on a minor inherent assumption that $Q_{N}$ satisfies $Q_{N}\succeq P^{*}$. A full discussion of this assumption is provided in Remark A.1 in Appendix A. Moreover, observe that since $P^{*}_{N}=Q_{N}\succeq P^{*}$, the monotonic convergence of (20) will yield
$Q_{N}=P^{*}_{N}\succeq P^{*}_{N-1}\succeq\cdots\succeq P^{*}_{0}\succeq P^{*},$
and hence
$\max_{t\in\{1,2,\dotsc,N\}}\{P^{*}_{t}\}=P^{*}_{N}=Q_{N}.$
Before presenting the formal algorithm, we provide an intuitive overview of its structure in Figure 1.
Freeze downstream policies fix $\widetilde{K}_{h+1},\dotsc,\widetilde{K}_{N-1}$ Build surrogate cost $\widetilde{J}_{h}(\cdot)$ function of the policy at $h$, with policies at [h+1,N1][h+1,N-1] fixed Initialize Estimate gradient $\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$ Update $K_{h,t+1}\!\leftarrow\!K_{h,t}-\alpha_{h,t}\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$ Output $\widetilde{K}_{h}\leftarrow K_{h,T_{h}}$ inner loop over $t=0,1,\dots,T_{h}$outer loop (receding horizon) over $h$: $N-1,\dots,1,0$
Figure 1. A brief overview of the nested-loop structure of Algorithm 1.
Figure 1 highlights the two nested loops: an outer loop stepping backward in $h$ (from N1N-1 to 0) and an inner loop iteratively refining the policy at each stage. At each $h$, the algorithm (i) freezes downstream policies, (ii) builds the surrogate cost $\widetilde{J}_{h}$, (iii) runs an inner loop of gradient-based updates with some step-size $\alpha_{h,t}$ until approximate convergence, and (iv) outputs $\widetilde{K}_{h}$, which is incorporated into the surrogate cost for step h1h-1.
We formalize this procedure in Algorithm 1 and briefly state what information it uses in practice:
@@
0: Horizon NN, max iterations $\{T_{h}\}$, stepsizes $\{\alpha_{h,t}\}$, variance parameter $\sigma^{2}$
1:for h=N1h=N-1 to 0 do
2:  Initialize $K_{h,0}$ arbitrarily (e.g., the convergent policy from the prev. iter. $K_{h+1,T_{h+1}}$ or 0).
@@
4:   Sample $\eta_{h,t}\sim\mathcal{N}(0,I_{m})$ and simulate a trajectory of the system (1) with
$u_{i}=\begin{cases}-K_{h,t}x_{0}+\sigma\eta_{h,t}&\text{: }i=0,\\ -\widetilde{K}_{h+i}x_{i}&\text{: }i\geq 1,\end{cases}$
and $x_{0}\sim\mathcal{D}$.
5:   Compute $Q_{h}(x_{0},u_{0})$ for said trajectory.
@@
$\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})\leftarrow-\frac{1}{\sigma}Q_{h}(x_{0},u_{h,t})\eta_{h,t}x_{0}^{\top}.$
7:   Update $K_{h,t+1}\leftarrow K_{h,t}-\alpha_{h,t}\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$.
@@
Information available to the algorithm. At each stage $h$, the updates require neither knowledge of the system matrices $(A,B)$ nor full trajectory information. Instead, the algorithm draws $\eta\sim\mathcal{N}(0,I)$, applies a perturbed action $u_{0}=-K_{h,t}x_{0}+\sigma\eta$ where $x_{0}\sim{\mathcal{D}}$, and then rolls out with the frozen downstream policies. It then queries a cost oracle for the finite-horizon cost-to-go $Q_{h}(x_{0},u_{0})$ of length NhN-h, which is used to form the gradient estimate for the inner updates. A detailed construction of $Q_{h}$ is deferred to Remark 4.1.
Note that in this section, we focus on the outer loop of Algorithm 1, analyzing the requirements it imposes on the convergence of the policy gradient method employed in the inner loop. The details of the policy gradient method will be discussed in the next section.
@@
$a:=\frac{\sigma_{\min}(Q)}{2},$
$\varphi:=\max_{t\in\{0,1,\dotsc,N-1\}}\|A-BK^{*}_{t}\|,$
$C_{1}:=\frac{\varphi\|B\|}{\lambda_{\min}(R)},$
$C_{2}:=2\varphi\|A\|\left(1+\frac{\|Q_{N}+aI\|\|B\|^{2}}{\lambda_{\min}(R)}\right),$
$C_{3}:=2\|R+B^{\top}(Q_{N}+aI)B\|.$
@@
(29) $\varsigma_{h,\varepsilon}:=\begin{cases}\displaystyle\min\!\left\{\begin{array}[]{l}\sqrt{\frac{a}{C_{3}}\frac{1}{N}},\;\sqrt{\frac{a}{C_{3}}\frac{a}{2eN\|Q_{N}\|}},\\[2.84526pt] \sqrt{\frac{a}{C_{3}}\frac{\varepsilon}{8eNC_{1}\|Q_{N}\|}},\;\sqrt{\frac{\varepsilon}{4C_{1}C_{3}}}\end{array}\right\}&h\geq 1,\\[5.69054pt] \displaystyle\frac{\varepsilon}{4}&h=0.\end{cases}$
We now present a key result, Theorem 3.2, on the accumulation of errors that constitutes an improvement over [22, Theorem 2] (corrected version of which is stated as Theorem 3.3 below); as the proof of Theorem 3.2 demonstrates, this improvement relies on a fundamentally different analysis.
@@
(30) $N=\frac{1}{2}\cdot\frac{\log\left(\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\varepsilon\cdot\lambda_{\min}(R)}\right)}{\log\left(\frac{1}{\|A_{K}^{*}\|_{*}}\right)}+1,$
where $Q_{N}\succeq P^{*}$, and suppose that Assumption 2.1 holds. Now assume that, for some $\varepsilon>0$, there exists a sequence of policies $\{\widetilde{K}_{h}\}_{h\in\{N-1,\dotsc,0\}}$ such that for all $h\in\{N-1,\dotsc,0\}$,
@@
where $\widetilde{K}^{*}_{h}$ is the optimal policy for the Linear Quadratic Regulator (LQR) problem from step $h$ to $N$, incorporating errors from previous iterations of Algorithm 1. Then, the proposed algorithm outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$. Moreover, for a sufficiently small $\varepsilon$ such that $\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\kappa_{P^{*}}^{1/2}\ \|B\|}$, $\widetilde{K}_{0}$ is stabilizing.
The proof of Theorem 3.2 is presented in Appendix B. A key component of our analysis is the contraction of the Riccati operator under the Riemannian distance, as established in Lemma 2.1. This allows us to demonstrate that the accumulated error remains linear in $N$, in contrast to the exponential growth in [22, Theorem 2].
Given this discrepancy, we revisit [22, Theorem 2] and present a revised version which accounts for some necessary, and non-trivial, modifications to make the statement accurate. For the latter reason, and the fact that this result does not rely on Assumption 2.1, we provide a complete proof in Appendix C.
@@
(31) $N=\frac{1}{2}\cdot\frac{\log\left(\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\varepsilon\cdot\lambda_{\min}(R)}\right)}{\log\left(\frac{1}{\|A_{K}^{*}\|_{*}}\right)}+1,$
where $Q_{N}\succeq P^{*}$. Now assume that, for some $\varepsilon>0$, there exists a sequence of policies $\{\widetilde{K}_{h}\}_{h\in\{N-1,\dotsc,0\}}$ such that
(32) $\|\widetilde{K}_{h}-\widetilde{K}^{*}_{h}\|\leq\begin{cases}\min\left\{\begin{array}[]{l}\sqrt{\frac{a}{C_{3}}},\sqrt{\frac{a}{C_{2}^{h-2}C_{3}}},\\ \frac{1}{2}\sqrt{\frac{\varepsilon}{C_{1}C_{2}^{h-2}C_{3}}}\end{array}\right\}&h\geq 2,\\[8.0pt] \min\left\{\begin{array}[]{l}\sqrt{\frac{a}{C_{3}}},\frac{1}{2}\sqrt{\frac{\varepsilon}{C_{1}C_{3}}}\end{array}\right\}&h=1,\\[8.0pt] \frac{\varepsilon}{4}&h=0,\end{cases}$
where $\widetilde{K}^{*}_{h}$ is the optimal policy for the Linear Quadratic Regulator (LQR) problem from step $h$ to $N$, incorporating errors from previous iterations of Algorithm 1. Then, the RHPG algorithm outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$.
@@ 4. The Inner Loop and Policy Gradient
We seek a way to estimate the gradient of the cost function $\widetilde{J}_{h}(K)$ with respect to $K$. To remedy, we propose:
(33) $\widehat{\nabla}\widetilde{J}_{h}(K):=-\frac{1}{\sigma^{2}}Q_{h}(x_{0},u_{0})(u_{0}+Kx_{0})x_{0}^{\top},$
where $x_{0}$ is sampled from $\mathcal{D}$, and then $u_{0}$ is chosen randomly from the Gaussian distribution $\mathcal{N}(-Kx_{0},\sigma^{2}I_{m})$ for some $\sigma>0$. Moreover, we rewrite $u_{0}\sim\mathcal{N}(-Kx_{0},\sigma^{2}I_{m})$ as
(34) $u_{0}=-Kx_{0}+\sigma\eta,$
where $\eta\sim\mathcal{N}(0,I_{m})$. Substituting (34) in (33) yields
(35) $\widehat{\nabla}\widetilde{J}_{h}(K)=-\frac{1}{\sigma}Q_{h}(x_{0},-Kx_{0}+\sigma\eta)\eta x_{0}^{\top}.$
@@
We now provide the following remark on the computation of $Q_{h}(x_{0},u_{0})$ for this gradient estimate:
Remark 4.1.
For any fixed $h$, the $Q$-function in (35) represents the finite-horizon cost-to-go of horizon length NhN-h, i.e., NhN-h stage costs followed by a terminal cost (indices in the notation can be perceived as shifts). To clarify, let us use the following notation
$u_{t}$ $:=\begin{cases}-Kx_{0}+\sigma\eta,&\text{: }t=0,\\ -\widetilde{K}_{h+t}\ x_{t},&\text{: }t\geq 1,\end{cases}$
where $(x_{t},u_{t})$ form the trajectory of the system, i.e., the dynamics follow
$x_{t+1}=Ax_{t}+Bu_{t},$
and $x_{0}\sim{\mathcal{D}}$. Now using the quadratic stage cost
$c^{(h)}_{t}:=\begin{cases}x_{t}^{\top}Q_{N}x_{t}&\text{: }t=N-h,\\ x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}&\text{: otherwise},\end{cases}$
we can write
$Q_{h}(x_{0},u_{0})=\sum_{t=0}^{N-h}c^{(h)}_{t}$.
Roadmap of the section. The core of our analysis in Section 4 relies on three key results: (i) unbiasedness of the gradient estimator (Proposition 4.1), (ii) a bound on the second-moment of the Frobenius norm of the estimate (Lemma 4.1), and (iii) regularity properties of the stage cost $\widetilde{J}_{h}$ (Lemma 4.2). Together, these yield a one-step descent recursion (Lemma 4.3), which leads to a probabilistic guarantee on the behavior of the optimality-gap (Proposition 4.2) and culminates in the main inner-loop convergence theorem (Theorem 4.1). Figure 2 summarizes these logical dependencies, with the three bolded results serving as the main pillars of the argument.
Proposition 4.1 (unbiasedness of $\widehat{\nabla}\widetilde{J}_{h}$) Lemma 4.1 (second-moment bound of $\widehat{\nabla}\widetilde{J}_{h}$) Lemma 4.2 (regularity properties of $\widetilde{J}_{h}$) Lemma 4.3 (one-step descent recursion) Proposition 4.2 (probabilistic optimality-gap guarantee) Theorem 4.1 (main inner-loop convergence guarantee)
Figure 2. Roadmap of technical results in Section 4.
We start by showing that the estimator is unbiased with respect to true gradient:
@@
Suppose $x_{0}\sim{\mathcal{D}}$ and $u_{0}$ is drawn from $\mathcal{N}(-Kx_{0},\sigma^{2}I_{m})$ as before. Then for any given choice of $K$, we have that
@@
${\mathbb{E}}[\widehat{\nabla}\widetilde{J}_{h}(K)]$ $={\mathbb{E}}_{x_{0}}\left[{\mathbb{E}}_{\eta}\left[\widehat{\nabla}\widetilde{J}_{h}(K)\big|x_{0}\right]\right]$
$\overset{\mathrm{(i)}}{=}{\mathbb{E}}_{x_{0}}\left[-\frac{1}{\sigma^{2}}{\mathbb{E}}_{\eta}\left[Q(x_{0},-Kx_{0}+\sigma\eta)(\sigma\eta)\big|x_{0}\right]x_{0}^{\top}\right]$
(37) $\overset{\mathrm{(ii)}}{=}{\mathbb{E}}_{x_{0}}\left[{\mathbb{E}}_{\eta}\left[-\nabla_{u}Q^{K}(x_{0},u)\bigg|_{u=-Kx_{0}+\sigma\eta}\big|x_{0}\right]x_{0}^{\top}\right],$
where (i) follows from $x_{0}^{\top}$ being determined when given $x_{0}$, and (ii) from Stein’s lemma [18]. Using (15), we compute
$\nabla_{u}Q_{h}(x_{0},u)$ $=\nabla_{u}\Bigg(x_{0}^{\top}Qx_{0}+u^{\top}Ru+(Ax_{0}+Bu)^{\top}\widetilde{P}_{h+1}(Ax_{0}+Bu)\Bigg)$
$=2Ru+2B^{\top}\widetilde{P}_{h+1}Bu+2B^{\top}\widetilde{P}_{h+1}Ax_{0},$
which evaluated at $u=-Kx_{0}+\sigma\eta$ yields
$\nabla_{u}Q_{h}(x_{0},u)\bigg|_{u=-Kx_{0}+\sigma\eta}=2\left((R+B^{\top}\widetilde{P}_{h+1}B)(-Kx_{0}+\sigma\eta)+B^{\top}\widetilde{P}_{h+1}Ax_{0}\right).$
@@
${\mathbb{E}}[\widehat{\nabla}\widetilde{J}_{h}(K)]$ $={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[2\left((R+B^{\top}\widetilde{P}_{h+1}B)K-B^{\top}\widetilde{P}_{h+1}A\right)x_{0}x_{0}^{\top}\right]$
$=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K-B^{\top}\widetilde{P}_{h+1}A\right){\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[x_{0}x_{0}^{\top}\right]$
@@
$\overset{\mathrm{(i)}}{=}\nabla\widetilde{J}_{h}(K),$
@@
We next provide a second-moment bound on the size of the estimator:
@@
Suppose $\|\widetilde{P}_{h+1}-P^{*}_{h+1}\|\leq a$. Then for any $K\in{\mathcal{G}}_{h}$, we have that
(39) ${\mathbb{E}}\left[\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}^{2}\right]\leq\xi_{h,3},$
where $\xi_{h,1},\xi_{h,2},\xi_{h,3}\in$ are given by
(40) $\xi_{h,1}$ $:=\Big(\|Q\|+2\|R\|\widetilde{C}_{h}^{2}+2(\|Q_{N}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h})\Big)C_{m}^{3/2},$
(41) $\xi_{h,2}$ $:=2\left(\|R\|+2(\|Q_{N}\|+a)\|B\|^{2}\right)C_{m}^{1/2},$
(42) $\xi_{h,3}$ $:=\frac{1}{\sigma^{2}}\xi_{h,1}^{2}m+2\xi_{h,1}\xi_{h,2}m(m+2)+\sigma^{2}\xi_{h,2}^{2}m(m+2)(m+4).$
@@
Using the Formulation of $\widehat{\nabla}\widetilde{J}_{h}(K)$ derived in (35), we have
(43) $\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}$ $=\|\frac{1}{\sigma}Q_{h}(x_{0},-Kx_{0}+\sigma\eta)\eta x_{0}^{\top}\|_{F}$
(44) $\leq\frac{1}{\sigma}Q_{h}(x_{0},-Kx_{0}+\sigma\eta)\|\eta\|\|x_{0}\|.$
@@
(46) $\widetilde{J}_{h}(K)$ $\geq{\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[x_{0}^{\top}(Q+K^{\top}RK)x_{0}\right]$
(47) $={\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[\operatorname{tr}\left((Q+K^{\top}RK)x_{0}x_{0}^{\top}\right)\right]$